### v3

@@ Abstract
Machine learning is playing an increasing role in hydrology, supplementing or replacing physics-based models. One notable example is the use of recurrent neural networks (RNNs) for forecasting streamflow given observed precipitation and geographic characteristics. Training of such a model over the continental United States (CONUS) demonstrated that a single set of model parameters can be used across independent catchments, and that RNNs can outperform physics-based models. In this work, we take a next step and study the performance of RNNs for river routing in land surface models (LSMs). Instead of observed precipitation, the LSM-RNN uses instantaneous runoff calculated from physics-based models as an input. We train the model with data from river basins spanning the globe and test it in streamflow hindcasts. The model demonstrates skill at generalization across basins (predicting streamflow in catchments not used in training) and across time (predicting streamflow during years not used in training). We compare the predictions from the LSM-RNN to an existing physics-based model calibrated with a similar dataset and find that the LSM-RNN outperforms the physics-based model. Our results show that RNNs are effective for global streamflow prediction from runoff inputs and motivate the development of complete routing models that can capture nested sub-basis connections.
@@ 1.3 Metrics
where $\hat{q}_{t}$ and $q_{t}$ are the predicted and observed streamflow at a basin outlet for time $t$ and $\bar{q}$ is the averaged observed streamflow over all valid times in the validation or test set. Note that here we omit the index $b$ because the expression is computed for each basin, and not for an ensemble of basins as we did in the expression of the loss function. In this expression, 1 is a perfect score ($\hat{q}_{t}=q_{t}$), and it gets worse (lower NSE) as the fraction of the mean squared error ($\sum_{t}(\hat{q}_{t}-q_{t})^{2}/n)$ normalized by the variance of the streamflow ($\sum_{t}(q_{t}-\bar{q})^{2}/n)$ increases, where $n$ indicates the number of observations for each basin. This normalization implies the NSE to lie in $(-\infty,1]$. Note that if a model is predicting only the mean flow at the outlet (i.e., $\hat{q}_{t}=\bar{q}$), we would have an NSE of 0. We will use this value as reference to evaluate good performances (NSE > 0) vs. bad performances (NSE < 0), as suggested in Knoben et al. (2019).
@@
In the equation, $r$ is the linear correlation coefficient between simulated and observed timeseries; $\alpha=\hat{\sigma}/\sigma$ is the variability ratio, given by the ratio between the standard deviation in simulations $\hat{\sigma}$ and the standard deviation in observations $\sigma$; $\beta=\hat{\mu}/\mu$ is the bias ratio, given by the ratio between the mean in simulations $\hat{\mu}$ and the mean in observations $\mu$. This score represents an explicit decomposition of the normalized mean squared error into the three components $r$, $\alpha$, and $\beta$. The score likewise lies in $(-\infty,1]$, with 1 being a perfect score. We can define a reference value based on the KGE for a model predicting only the mean flow. In this case, we would have no correlation ($r=0$), no variability ratio ($\alpha=0$), but a perfect bias ratio ($\beta=1$), which gives a reference KGE of $0.411-\sqrt{2}\approx-$ 0.41. We will use this reference value as a parameter to evaluate a good performance (KGE > $1-\sqrt{2}$) vs. a bad performance (KGE < $1-\sqrt{2}$).
@@ 2 Results
We present a series of experiments carried out to quantify the performance of the model. We compare the behavior of a model driven with precipitation to that of one driven with runoff, and we compare the behavior of a model trained and tested in the USA with one trained and tested globally. To assess generalizability across basins and times, we experiment with different training/testing/validation splits. The model’s training time varied with the datasets and the longest run took a few hours with one V100 GPU. Furthermore, we compare the performance of the model against the physics-based model LISFLOOD (Van Der Knijff et al., 2010), provided by GloFAS v4.0 reanalysis data, which is publicy available in the Copernicus Climate Data Store (https://confluence.ecmwf.int/display/CEMS/GloFAS+v4.0, last accessed 4 September 2024). The hindcasts generated by both models are presented under various conditions, including those where both models produce poor hindcasts, those where the hindcasts are close to the median score, and those where each model demonstrates good performance. An additional point is made regarding possible comparisons with the models presented here and other LSTM models in the literature. Finally, we analyze the performance by continent and other attributes. An analysis by HydroSHEDS levels is provided in Appendix C. All models shown in this section are based on the same LSTM architecture, which is not strictly mass conserving. An analysis of mass conservation for our LSTM can be found in Appendix D.
@@ 2.1.3 Summary of results
Table 2 highlights some important additional points. The time-split configuration trained on observed USA precipitation data has the smallest fraction of bad performances among the configurations trained and tested in the USA, suggesting that it suffers less from outliers and that there is room for improvement in runoff modeling. We also highlight that MeanNSE>0 is a good quantitative number to summarize the overall behavior of the curves in the CDF in Figure 4.
@@ 2.2.1 Comparison with a physics-based model
In a second series of experiments, we compare the LSTM model with the LISFLOOD model, provided by GloFAS v4.0. This choice was made because LISFLOOD explicitly calculates runoff and performs routing of river water from reanalysis data from ERA5 (Hersbach et al., 2020). Moreover, the comparison between the LSTM and LISFLOOD was straightforward to carry out without additional calibrations, as the LISFLOOD streamflow predictions are provided with the reanalysis data. That said, there are some important differences between our model and training procedure and those of GloFAS: (i) GloFAS uses different forcings from ERA5, including precipitation, to calculate runoff internally, whereas the LSTM uses runoff variables computed by the HTESSEL model (Balsamo et al., 2009) in ERA5-Land (ii) LISFLOOD is a complete river routing model, whereas the LSTM models basins independently; (iii) the objective function for GloFAS is based on the KGE metric whereas the LSTM was trained to optimize the NSE; and (iv) GloFAS is calibrated with a different global set of gauges.
@@
For the gauges we use, the LSTM shows an overall better accuracy compared with GloFAS in both basin- and time- split experiments. For the temporal split, the LSTM has a median score of 0.64 on NSE and 0.72 on KGE, whereas GloFAS displays a median score of 0.53 on NSE and 0.71 on KGE. For the basin split, the LSTM has a median score of 0.33 on NSE and 0.41 on KGE, whereas GloFAS displays a median score of 0.30 on NSE and 0.40 on KGE. Observe that despite being optimized for NSE, the LSTM has superior scores on both basin- and time-split to GloFAS on KGE. Moreover, the LSTM predictions result in fewer basins with performance worse than the baseline in all experiments. We show other metrics in Table 3.
@@ 2.2.2 Comparison with other LSTM models
In this section, we present some simulated timeseries along with observations for various values of NSE and KGE to get a visual sense of performance. For four different basins, Figure 6 shows the predictions of the LSTM model in the global time-split configuration, the GloFAS reanalysis, and the observed streamflow at the corresponding GRDC gauge.
@@ 3 Discussion and conclusion
Long short-term memory models have been shown to be the state of the art for modeling streamflow in hydrological systems (Kratzert et al., 2019b), but most studies using these models have so far been restricted to specific regions, and they have focused on representing the entire land hydrological system as a single entity (Kratzert et al., 2019b; Koch and Schneider, 2022). However, for integration into the land surface models used in climate modeling, it is crucial for river models to demonstrate proficiency in generalizing across basins and time scales globally, using surface and subsurface runoff data, rather than precipitation data. This necessity arises because traditional river models in land surface models route water, but leave the modeling of snow, soil, and vegetation hydrological processes to other model components. In this study, we have taken concrete steps toward developing an ML model that can substitute traditional river routing models within LSMs.
@@
Additionally, our analysis revealed a correlation between the model’s performance and the aridity index. While streamflow in arid basins can be modeled well by the LSTM, it is also true that all basins with a poor NSE have a lower aridity index. This suggests that drier regions pose challenges for the LSTM model, but that other basin features may affect performance as well. The dependence on aridity could be due to the model’s difficulty in capturing short-term precipitation events that trigger streamflow peaks in these areas, as seen in Figure 6a (Feng et al., 2020). It also underperformed in regions likely underrepresented in the training data. Despite these challenges, the model’s time generalizability was consistent across different basin sizes, even if better basin generalizability was observed for bigger basins, as discussed in Appendix C. We highlight that the model doesn’t need to be re-calibrated for the evaluation of different basin levels, which correspond to different spatial resolutions of LSMs.
@@
In contrast, the physics-based LISFLOOD model (GloFAS) is more similar to our model in that it uses modeled runoff (calculated internally) as a dynamic input to perform streamflow routing. We have compared the models on a subset of the gauges used to calibrate GloFAS (time-split) and on a subset of gauges not used for calibration (basin-split). The LSTM model showed more accurate simulations in both scenarios, but note that there are still some differences in experimental design between the two. Unfortunately, a recalibration of the model using our data and experimental design is beyond the scope of this work – previous works have reported calibration of GloFAS that took months in computational time (Alfieri et al., 2020).
In order to integrate our model in with an LSM, it must be extended to channel routing and it must respect physical priniciples like mass conservation. To our knowledge, no article has demonstrated a machine learning-based approach for global routing simulations, although some approaches show how to account for inter-basin connections in a restricted region (Moshe et al., 2020). On the other hand, physical models such as LISFLOOD model routing between basins and hence are currently the standard for LSMs. One may interpret this work as the hillslope routing component of a complete routing model. In that case, two possible ways to extend this to route water between basins are to use a physics-based channel routing model or to use similar neural networks to additionally perform river channel routing. For the latter, designing training strategies to mitigate error accumulation across connected basins may be necessary. Moreover, incorporating architectural adjustments to ensure mass conservation (e.g., Hoedt et al., 2021), is a necessary step for coupling with climate models, potentially enhancing long-term simulation accuracy. In conclusion, we have demonstrated that machine learning based river models exhibit basin and time generalizability, a requirement for use in a global climate model. Our study presents the first step toward using these models in the LSM component of climate models and our results motivate further research to extend these models for comprehensive river routing.
@@ Acknowledgements.
This research was supported by the generosity of Eric and Wendy Schmidt by recommendation of the Schmidt Futures program. We thank Frederik Kratzert for a helpful early conversation, as well as the other maintainers of neuralhydrology, for the clear and helpful code. We thank Svea Bertolatus, Stefania Grimaldi and the GloFAS team for providing the set of gauges from GRDC used to calibrate the model that was used to download the discharge reanalysis data, as well as helpful clarifications about GloFAS versions. ML acknowledges Akshay Sridhar, who assisted with technical parts of the code, data handling and insightful conversations, as well as Thomas Dubos for helping this research project to take place. All numerical calculations and model training used for this manuscript were performed with the help of Caltech’s Resnick High Performance Computing Center.

### v4

@@ Abstract
Machine learning is playing an increasing role in hydrology, supplementing or replacing physics-based models. One notable example is the use of recurrent neural networks (RNNs) for forecasting streamflow given observed precipitation and geographic characteristics. Training of such a model over the continental United States (CONUS) demonstrated that a single set of model parameters can be used across independent catchments, and that RNNs can outperform physics-based models. In this work, we take a next step and study the performance of RNNs for river routing in land surface models (LSMs). Instead of observed precipitation, the LSM-RNN uses instantaneous runoff calculated from physics-based models as an input. We train the model with data from river basins spanning the globe and test using historical streamflow measurements. The model demonstrates skill at generalization across basins (predicting streamflow in catchments not used in training) and across time (predicting streamflow during years not used in training). We compare the predictions from the LSM-RNN to an existing physics-based model calibrated with a similar dataset and find that the LSM-RNN outperforms the physics-based model. Our results show that RNNs are effective for global streamflow prediction from runoff inputs and motivate the development of complete routing models that can capture nested sub-basis connections.
@@ 1.3 Metrics
where $\hat{q}_{t}$ and $q_{t}$ are the predicted and observed streamflow at a basin outlet for time $t$ and $\bar{q}$ is the averaged observed streamflow over all valid times in the validation or test set. Note that here we omit the index $b$ because the expression is computed for each basin, and not for an ensemble of basins as we did in the expression of the loss function. In this expression, 1 is a perfect score ($\hat{q}_{t}=q_{t}$), and it gets worse (lower NSE) as the fraction of the mean squared error ($\sum_{t}(\hat{q}_{t}-q_{t})^{2}/n)$ normalized by the variance of the streamflow ($\sum_{t}(q_{t}-\bar{q})^{2}/n)$ increases, where $n$ indicates the number of observations for each basin. This normalization implies the NSE to lie in $(-\infty,1]$. Note that if a model is predicting only the mean flow at the outlet (i.e., $\hat{q}_{t}=\bar{q}$), we would have an NSE of 0. We will use this value as reference to evaluate performances above a naive mean flow baseline (NSE > 0) vs. performances below a naive mean flow baseline (NSE < 0), as suggested in Knoben et al. (2019).
@@
In the equation, $r$ is the linear correlation coefficient between simulated and observed timeseries; $\alpha=\hat{\sigma}/\sigma$ is the variability ratio, given by the ratio between the standard deviation in simulations $\hat{\sigma}$ and the standard deviation in observations $\sigma$; $\beta=\hat{\mu}/\mu$ is the bias ratio, given by the ratio between the mean in simulations $\hat{\mu}$ and the mean in observations $\mu$. This score represents an explicit decomposition of the normalized mean squared error into the three components $r$, $\alpha$, and $\beta$. The score likewise lies in $(-\infty,1]$, with 1 being a perfect score. We can define a reference value based on the KGE for a model predicting only the mean flow. In this case, we would have no correlation ($r=0$), no variability ratio ($\alpha=0$), but a perfect bias ratio ($\beta=1$), which gives a reference KGE of $0.411-\sqrt{2}\approx-$ 0.41. We will use this reference value as a parameter to evaluate performances above a naive mean flow baseline (KGE > $1-\sqrt{2}$) vs. performances below a naive mean flow baseline (KGE < $1-\sqrt{2}$).
@@ 2 Results
We present a series of experiments carried out to quantify the performance of the model. We compare the behavior of a model driven with precipitation to that of one driven with runoff, and we compare the behavior of a model trained and tested in the USA with one trained and tested globally. To assess generalizability across basins and times, we experiment with different training/testing/validation splits. The model’s training time varied with the datasets and the longest run took a few hours with one V100 GPU. Furthermore, we compare the performance of the model against the physics-based model LISFLOOD (Van Der Knijff et al., 2010), provided by GloFAS v4.0 reanalysis data, which is publicy available in the Copernicus Climate Data Store (https://confluence.ecmwf.int/display/CEMS/GloFAS+v4.0, last accessed 4 September 2024). The simulations generated by both models are presented under various conditions, including those where both models produce poor scores, those where the simulations are close to the median score, and those where each model demonstrates good performance relative to the median score of each model. An additional point is made regarding possible comparisons with the models presented here and other LSTM models in the literature. Finally, we analyze the performance by continent and other attributes. An analysis by HydroSHEDS levels is provided in Appendix C. All models shown in this section are based on the same LSTM architecture, which is not strictly mass conserving. An analysis of mass conservation for our LSTM can be found in Appendix D.
@@ 2.1.3 Summary of results
Table 2 highlights some important additional points. The time-split configuration trained on observed USA precipitation data has the smallest fraction of performances with scores worse than the naive mean-flow baseline model among the configurations trained and tested in the USA, suggesting that it suffers less from outliers and that there is room for improvement in runoff modeling. We also highlight that MeanNSE>0 is a good quantitative number to summarize the overall behavior of the curves in the CDF in Figure 4.
@@ 2.2.1 Comparison with a physics-based model
In a second series of experiments, we compare the LSTM model with the LISFLOOD model, provided by GloFAS v4.0. This choice was made because LISFLOOD explicitly calculates runoff and performs routing of river water from reanalysis data from ERA5 (Hersbach et al., 2020). Moreover, the comparison between the LSTM and LISFLOOD was straightforward to carry out without additional calibrations, as the LISFLOOD streamflow predictions are provided with the reanalysis data. That said, there are some important differences between our model and training procedure and those of GloFAS: (i) GloFAS uses different forcings from ERA5, including precipitation, to calculate runoff internally, whereas the LSTM uses runoff variables computed by the HTESSEL model (Balsamo et al., 2009) in ERA5-Land (ii) LISFLOOD is a complete river routing model, whereas the LSTM models basins independently; (iii) the objective function for GloFAS is based on the modified KGE metric whereas the LSTM was trained to optimize the NSE; and (iv) GloFAS is calibrated with a different global set of gauges.
@@
For the gauges we use, the LSTM shows an overall better accuracy compared with GloFAS in both basin- and time- split experiments. For the temporal split, the LSTM has a median score of 0.64 on NSE and 0.72 on KGE, whereas GloFAS displays a median score of 0.53 on NSE and 0.71 on KGE. For the basin split, the LSTM has a median score of 0.33 on NSE and 0.41 on KGE, whereas GloFAS displays a median score of 0.30 on NSE and 0.40 on KGE. Observe that despite being optimized for NSE, the LSTM has similar scores on both basin- and time-split to GloFAS on KGE. Moreover, the LSTM predictions result in fewer basins with performance worse than the baseline in all experiments. We show other metrics in Table 3.
@@ 2.2.2 Comparison with other LSTM models
In this section, we present simulated timeseries along with observations for various values of NSE and KGE to get a visual sense of performance. For four different basins, Figure 6 shows the predictions of the LSTM model in the global time-split configuration, the GloFAS reanalysis, and the observed streamflow at the corresponding GRDC gauge.
@@ 3 Discussion and conclusion
Long short-term memory models have been shown to be the state of the art for modeling streamflow in hydrological systems (Kratzert et al., 2019b), but most studies using these models have so far been restricted to specific regions, and they have focused on representing the entire land hydrological system as a single entity (Kratzert et al., 2019b; Koch and Schneider, 2022). However, for integration into the land surface models used in weather forecasting and climate modeling, it is crucial for river models to demonstrate proficiency in generalizing across basins and time scales globally, using surface and subsurface runoff data, rather than precipitation data. This necessity arises because traditional river models in land surface models route water, but leave the modeling of snow, soil, and vegetation hydrological processes to other model components. In this study, we have taken concrete steps toward developing an ML model that can substitute traditional river routing models within LSMs.
@@
Additionally, our analysis revealed a correlation between the model’s performance and the aridity index. While streamflow in arid basins can be modeled well by the LSTM, it is also true that all basins with a poor NSE have a lower aridity index. This suggests that drier regions pose challenges for the LSTM model, but that other basin features may affect performance as well. The dependence on aridity could be due to the model’s difficulty in capturing short-term precipitation events that trigger streamflow peaks in these areas, as seen in Figure 6a (Feng et al., 2020). Moreover, the ERA5-Land reanalysis is driven with precipitation from ERA5, known to have biases in the tropics (Lavers et al., 2022), which could lead to biases in the runoff of ERA5-Land in these regions. As a consequence, our river model could be learning to correct these biases in addition to routing water. This is a possible pitfall for machine learning models trained with model output and not with observed data. The LSTM model also underperformed in regions likely underrepresented in the training data. Despite these challenges, the model’s time generalizability was consistent across different basin sizes, even if better basin generalizability was observed for bigger basins, as discussed in Appendix C. We highlight that the model does not need to be re-calibrated for the evaluation of different basin levels, which correspond to different spatial resolutions of LSMs.
@@
In contrast, the physics-based LISFLOOD model (GloFAS) is more similar to our model in that it uses modeled runoff (calculated internally) as a dynamic input to perform streamflow routing. We have compared the models on a subset of the gauges used to calibrate GloFAS (time-split) and on a subset of gauges not used for calibration (basin-split). The LSTM model showed more accurate simulations in both scenarios, but note that there are still some differences in experimental design between the two. Unfortunately, a recalibration of the model using our data and experimental design is beyond the scope of this work – the reported calibration time of GloFAS is of the order of weeks (GloFAS team members, private communication).
In order to integrate our model with an LSM, it must be extended to channel routing and it must respect physical priniciples like mass conservation. To our knowledge, no article has demonstrated a machine learning-based approach for global routing simulations, although some approaches show how to account for inter-basin connections in a restricted region (Moshe et al., 2020). On the other hand, physical models such as LISFLOOD model routing between basins and hence are currently the standard for LSMs. One may interpret this work as the hillslope routing component of a complete routing model. In that case, two possible ways to extend this to route water between basins are to use a physics-based channel routing model or to use similar neural networks to additionally perform river channel routing. For the latter, designing training strategies to mitigate error accumulation across connected basins may be necessary. Moreover, incorporating architectural adjustments to ensure mass conservation (e.g., Hoedt et al., 2021), is a necessary step for coupling with climate models, potentially enhancing long-term simulation accuracy, although this might not be necessary for integration into LSMs used by weather forecasting models. In conclusion, we have demonstrated that machine learning based river models exhibit basin and time generalizability, a requirement for use in a global climate model. Our study presents the first step toward using these models in the LSM component of climate models as well as for short, medium and long range hydrological forecasting using runoff from LSMs within weather forecasting models. The results presented here motivate further research to extend these models for comprehensive river routing.
@@ Acknowledgements.
This research was supported by the generosity of Eric and Wendy Schmidt by recommendation of the Schmidt Futures program. We thank Frederik Kratzert for a helpful early conversation, as well as the other maintainers of neuralhydrology, for the clear and helpful code. We thank the CESM GloFAS team for providing the set of gauges from GRDC used to calibrate the model that was used to download the discharge reanalysis data, as well as helpful clarifications about GloFAS versions. ML acknowledges Akshay Sridhar, who assisted with technical parts of the code, data handling and insightful conversations, as well as Thomas Dubos for helping this research project to take place. All numerical calculations and model training used for this manuscript were performed with the help of Caltech’s Resnick High Performance Computing Center.